OpenStudy (anonymous):

consider the following quadratic equation: x^2-7x+1=9 using the standard form ax^2+bx+c=0 of the quadratic equation above, factor the left hand side equation into two linear factors

jhonyy9 (jhonyy9):

x*2 -7x +1 =9 x*2 - 7x -8 =0 (x-8)(x+1)=0 x_1=1 x_2=-8

OpenStudy (anonymous):

\[(x-8)(x+1)\]